Which equations is undefined
A) y=4
B) y=-2x
C) x=4
D) y=x+4

Found 2 solutions by Nate, rapaljer:
Answer by Nate(3500)

Vertical lines are undefined:
x+=+4
take two points: (4,0) and (4,2)
M = (y2 - y1)/(x2 - x1) = (0 - 2)/(4 - 4) = -2/0 undefined
(C) is your answer.

Answer by rapaljer(4671)

Do you mean which equation has a SLOPE that is undefined? If so, then this would be the vertical line, which is C) x= 4.

R^2 at SCC